what is the function for all these number
-6,-5
-4,-4
-2,-3
0,-2
2,-1
4,0
6,1
8,2
they all have to have the same function

Answer by martinr(6)

You need to find the slope and the y-intercept for this problem.
The line is linear, if you were to graph the points, all the points will be on the same line (Straight line).
Slope is the difference in y's over the difference is x's.
I will use (-6, -5) (-4, -4)
m = -4 - -5 = 1
-4 - -6 = 2
therfore the slope is m = 1/2 or .5.
Next find the y-intecept.
First choose one of the points (-6, -5). (Makes no difference).
y = mx + b, substitute for x, y and m.
-5 = (1/2)(-6) + b
-5 = -3 + b (add three to both sides)
b = -2
Therefore the equation of the line is y = (1/2)x + -2.
You can use the same process for other points and if you really have a lot of time, you can use every combination of points and notice that they all have the same equation.
Good luck.